OpenStudy (anonymous):

You have a bag containing 26 tiles, each imprinted with one letter of the alphabet. If you pick out 5 tiles, what is the probability that they are a, e, i, o, and u in that order?

OpenStudy (anonymous):

\[\frac{ 1 }{ 26 } \times \frac{ 1 }{ 25 } \times \frac{ 1 }{ 24 } \times \frac{ 1 }{ 23 } \times \frac{ 1 }{ 22 }\]

OpenStudy (anonymous):

what is the outcome

OpenStudy (anonymous):

or how do i make a fraction on the calculator

OpenStudy (anonymous):

@Doclav

OpenStudy (astriddawnrose):

the answer is D future ppl or 1/7893600